# Meta-Learning an In-Context Transformer Model of Human Higher Visual Cortex

### v1

@@ A.8 Voxelwise explained-variance evaluation in BOLD5000 for more subjects and different backbones
A.9 Dimensional reduction of predicted response function weights on more subjects
@@
Figure S.14: Dimensional reduction of predicted response weights for subject S1-S4 under CLIP backbone. The cortical maps show color-coded mappings that align well with functionally-defined regions: body and face regions (EBA and FFA/aTL-faces), place regions (RSC/OPA/PPA), and food regions (in red).
Figure S.15: Dimensional reduction of predicted response weights for subject S5-S8 under CLIP backbone. The cortical maps show color-coded mappings that align well with functionally-defined regions: body and face regions (EBA and FFA/aTL-faces), place regions (RSC/OPA/PPA), and food regions (in red).
A.10 Predicting cortical responses from natural language prompts on more subjects
@@
A.11 Voxelwise prompt classification accuracy for more subjects
In this section, we further quantify the semantic specificity of BraInCoRL’s voxelwise predictions on subject 3-8. we compute, for each subject and each category-selective ROI, the fraction of voxels whose peak predicted activation corresponded to the semantic category named by the text prompt.

### v2

@@ A.10 Voxelwise explained-variance evaluation in BOLD5000 for more subjects and different backbones
A.11 Dimensional reduction of predicted response function weights on more subjects
@@
Figure S.14: Dimensional reduction of predicted response weights for subject S1-S4 under CLIP backbone. The cortical maps show color‐coded mappings that align well with functionally‐defined regions: body and face regions (EBA and FFA/aTL‐faces), place regions (RSC/OPA/PPA), and food regions (in red).
Figure S.15: Dimensional reduction of predicted response weights for subject S5-S8 under CLIP backbone. The cortical maps show color‐coded mappings that align well with functionally‐defined regions: body and face regions (EBA and FFA/aTL‐faces), place regions (RSC/OPA/PPA), and food regions (in red).
A.12 Predicting cortical responses from natural language prompts on more subjects
@@
A.13 Voxelwise prompt classification accuracy for more subjects
In this section, we further quantify the semantic specificity of BraInCoRL’s voxelwise predictions on subject 3-8. we compute, for each subject and each category‐selective ROI, the fraction of voxels whose peak predicted activation corresponded to the semantic category named by the text prompt.
@@ A.13 Voxelwise prompt classification accuracy for more subjects
A.14 Additional evaluation of BraInCoRL on NSD dataset
In this section, we provide two more evaluation metrics, namely Pearson $R$, and Spearman’s rank correlation coefficient (Spearman’s $\rho$) for NSD dataset on Subject 1. In this case, the BraInCoRL model has not been trained or finetuned on Subject 1, while the Fully Trained model is trained on 9,000 images from this subject.
Figure S.23: Voxel-wise Pearson $R$ for BraInCoRL, within-subject ridge regression baseline and fully-trained reference model (NSD dataset, CLIP backbone, Subject 1, higher is better).
Figure S.24: Voxel-wise Spearman’s $\rho$ for BraInCoRL, within-subject ridge regression baseline and fully-trained reference model (NSD dataset, CLIP backbone, Subject 1, higher is better).
A.15 Additional evaluation of BraInCoRL on BOLD5000 dataset
In this section, we provide two more evaluation metrics, namely explained variance and Spearman’s rank correlation coefficient (Spearman’s $\rho$) for BOLD5000 dataset on Subject CSI1.
Figure S.25: Voxel-wise explained variance for BraInCoRLand within-subject ridge regression baseline (BOLD5000 dataset, CLIP backbone, Subject CSI1, higher is better).
Figure S.26: Voxel-wise Spearman’s $\rho$ for BraInCoRLand within-subject ridge regression baseline (BOLD5000 dataset, CLIP backbone, Subject CSI1, higher is better).
A.16 Evaluation of each training stage’s contribution
In this section, we present the voxelwise explained variance for BraInCoRLtrained until different training stages. The results show that the progression from synthetic foundation → context flexibility → biological adaptation ensures that each fundamental challenge, namely response function coverage, variable context handling, and biological realism, is systematically addressed in the optimal order.
Figure S.27: Voxelwise explained variance for BraInCoRLwith CLIP backbone for NSD Subject 1 on different training stages, compared with ridge baseline and fully trained reference model (higher is better).
Figure S.28: Voxelwise explained variance for BraInCoRLwith CLIP backbone for NSD Subject 2 on different training stages, compared with ridge baseline and fully trained reference model (higher is better).
A.17 Performance of BraInCoRL conditioned on the full 9000-image set
In this section, we evaluate the performance of BraInCoRL with 9,000 in-context samples, for the four NSD subjects (S1, S2, S5, S7) focues by our main paper. The performance difference is less than 1% across all subjects when compared to a fully trained model (which is fit to converge on each subject’s entire 9,000-image training set using gradient descent over multiple epochs). This means BraInCoRL achieves 94-99% of the fully trained model’s performance.
Table S.16: Voxel-wise explained variance of BraInCoRL with the CLIP backbone compared with the fully trained reference model. The difference variance explained is less than 1%.
Subject
Method S1 S2 S5 S7
Fully Trained 0.1765 0.1882 0.2310 0.1554
BrainCoRL 0.1667 0.1817 0.2225 0.1541
A.18 Evaluation on the choice of loss function during training
In this section, we conducted an ablation study on the choice of different loss functions during the finetuning stage optimization and evaluated the model performance.
Our experimental results show that MSE and L1 losses achieve similar performance across all context sizes, with minimal differences. This suggests that both metrics are equally effective for capturing voxelwise neural response patterns.
In addition, the hybrid loss of $0.5\times\text{MSE loss}+0.5\times(1-\text{cosine similarity})$ underperforms by approximately 2-4% compared to MSE/L1. We argue this is because although cosine similarity captures directional relationships between predicted and true responses, this additional constraint limit the model’s ability to accurately predict response magnitudes.
Figure S.29: Voxel-wise explained variance of different training losses with the CLIP backbone for Subject 1 (higher is better).
A.19 Ablation on logit scaling
In this section, we provide an additional ablation study to evaluate the effect of the logit-scaling, where we report the Voxel-wise explained variance of our BrainCoRL model and the model with the exact same structure but without logit scaling. It is shown that the logit scaling significantly boosts the model performance and generizability of various in-context support set sizes.
Figure S.30: Voxel-wise explained variance for BrainCoRL with CLIP backbone for Subj 1, compared to the same model architecture but without logit scaling (higher is better).